OpenStudy (commonlizer):

A system of equations is shown below: y = 8x − 2 y = 9x − 7 What is the solution to the system of equations? A (−5, 38) B (−5, −38) C (5, 38) D (5, −38)

OpenStudy (commonlizer):

just checking this my answer, my answer is B am I right

rebeccaxhawaii (rebeccaxhawaii):

okay so y=8x-2 so plug it into the second equation for y 8x-2=9x-7

OpenStudy (commonlizer):

8-2 = 6 and 9-7 = 2

rebeccaxhawaii (rebeccaxhawaii):

no you need to add like terms so 8x-9x=-7+2 whats 8x-9x=? and -7+2=?

OpenStudy (commonlizer):

oh so 8-9 = 1 and -7+2 = -9

rebeccaxhawaii (rebeccaxhawaii):

no you have 8 oranges you owe 9 oranges. so you give your 8 how many do you still have to give?

OpenStudy (commonlizer):

One more right?

rebeccaxhawaii (rebeccaxhawaii):

yes so -1

rebeccaxhawaii (rebeccaxhawaii):

8-9=-1

rebeccaxhawaii (rebeccaxhawaii):

if you owe 7 oranges and you only have 2, so you give your two how many do you still owe?

OpenStudy (commonlizer):

5 Oranges

rebeccaxhawaii (rebeccaxhawaii):

correct so -7+2=-5

rebeccaxhawaii (rebeccaxhawaii):

-x=-5 divide both by a negative it will become x=5

OpenStudy (commonlizer):

okay so we have -1 and x=5

OpenStudy (commonlizer):

but this is making no sense to me because all of the answers are 5 and 38

rebeccaxhawaii (rebeccaxhawaii):

nope we arent done all we know so far is that x=5

OpenStudy (commonlizer):

oh okay let's continue

rebeccaxhawaii (rebeccaxhawaii):

so plug in 5 for x in one of the equations y=8x-2 y=8(5)-2

OpenStudy (commonlizer):

8 x 5 = 40 - 2 = 38 right?

rebeccaxhawaii (rebeccaxhawaii):

correct so x=5 and y=38 (x, y) (5,38)

OpenStudy (commonlizer):

thank you for your help once again

rebeccaxhawaii (rebeccaxhawaii):

no problem c: